OpenStudy (anonymous):

Emergency Looking for an angel to save me before I fail math class!! Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (triciaal):

if c is a root of an equation then (x-c) is a factor when you divide the function by (x-c) the remainder = 0 difference means subtract squares means power 2

OpenStudy (anonymous):

I'm not following you, I'm very confused because the problem is asking to create 4 different functions but I'm stuck on that

OpenStudy (triciaal):

g(x) sum means to add h(x) trinomial means it has 3 terms perfect square means the root is an integer

OpenStudy (anonymous):

oh ok, could you give me an example of those please

OpenStudy (triciaal):

each situation is a different function f(x) = a^2 - b^2 = (a - b)(a +b)

OpenStudy (anonymous):

ok tyvm :)

OpenStudy (triciaal):

(x+5)^2 = x^2 + 10x + 25

OpenStudy (triciaal):

you are welcome

OpenStudy (anonymous):

would that right there be a perfect square trinomial example :)?

OpenStudy (triciaal):

yes 5^2 = 25 3 terms in the function

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

could you maybe give me an example of g(x) and j(x) now :D?

OpenStudy (triciaal):

compare f(x) difference means to subtract g(x) sum means to add you can do that

OpenStudy (triciaal):

j(x) to have a common factor whatever the function is, each coefficient can be divided by the GCF

OpenStudy (triciaal):